Find the possible values of k for which the quadratic equation 9x^2- 2(k+7)x + k=0 has equal value.
Answer by mananth(16946)


9x^2- 2(k+7)x + k=0
compare equation with ax^2+bx+c
a= 9 , b=-2(k+7) , c = k
if roots are equal
b^2-4ac =0
But the roots are not real